# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ V-A Massive MIMO
MMSE-based channel estimation is commonly used in MIMO systems to estimate the channel response in noisy environments. It minimizes the mean squared error between the actual channel and its estimate using the second-order statistics of the channel and noise. In massive MIMO systems with FDD, reducing CSI feedback overhead is challenging. A method in [62] combines deep learning and superposition coding to address this. Downlink CSI is encoded, superimposed on uplink user data, and sent to the base station. A multi-task neural network unfolds the MMSE criterion to reduce interference and reconstruct the CSI efficiently.
The iterative shrinkage-thresholding algorithm (ISTA) is commonly used for sparse channel estimation, enabling efficient CSI feedback. Each ISTA iteration includes a linear operation followed by a nonlinear soft-thresholding step, similar to the ReLU activation function [16]. In [63], Chen et al. proposed a trainable deep learning architecture for sparse recovery problems using an adaptive depth approach. Instead of a fixed number of layers, the method dynamically adjusts the depth based on the problem’s complexity. Highly sparse channels use fewer layers, while more complex channels require additional layers. This adaptive approach improves accuracy and lowers computational costs compared to traditional methods and fixed-depth networks. Results in [63] demonstrate that the method reduces reconstruction error and speeds up channel estimation, making it more efficient.
A scenario-adaptive CSI compression framework is proposed in [168], requiring minimal channel measurements for training. The authors introduced a CSI feedback architecture designed for efficient encoder updates in dynamic wireless environments, which is essential for downlink precoding in massive MIMO systems. The encoding network focuses on improving the robustness and efficiency of CSI compression. It employs a deep unfolding-based feedback network to iteratively refine the feedback process and enhance feature extraction. Similarly, the decoding network uses a deep unfolding structure inspired by the ISTA algorithm. This iterative approach improves CSI recovery accuracy, making the process more efficient and adaptable to changing conditions.
In massive MIMO systems, where the channel matrix is often sparse, the majorization-minimization (MM) approach is effective for addressing sparse channel estimation [169]. In [64], the authors address the critical challenge of downlink channel estimation in massive MIMO, essential for efficient communication. Traditional methods often rely on angular-domain channel sparsity to reduce pilot overhead but struggle in dynamic environments. To overcome this, the authors propose an online two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm. This method leverages channel sparsity and angle reciprocity between uplink and downlink transmissions to adapt to dynamic scenarios and significantly reduce pilot overhead. The two-timescale strategy includes learning a long-term dictionary to capture spatial channel characteristics and frequently estimating short-term sparse channel vectors. The short-term channel estimation subproblems are solved using an unfolded MM algorithm, while the long-term dictionary optimization uses standard convex methods. Deep unfolding is employed to jointly optimize the dictionary learning and channel estimation process for both uplink and downlink. Simulations in [64] show that this approach reduces pilot overhead and outperforms existing methods in static and dynamic environments, especially in scenarios with mobile users.
The matching pursuit algorithm is a greedy iterative method often applied in sparse channel scenarios like massive MIMO. It identifies channel components (basis functions) by matching them to the residual, which is the difference between the observed signal and the signal generated by the current channel estimate. Yassine et al. [65] introduced a deep unfolded algorithm, called mpNet, based on the classical matching pursuit algorithm. This network estimates the multiuser massive MIMO channel in an unsupervised manner, initializing with a dictionary of imperfect steering vectors. The system operates in time division duplex (TDD) mode to leverage channel reciprocity. The network dynamically adjusts the number of layers based on incoming data, ensuring efficient performance in realistic environments.
The basis pursuit algorithm is a widely used method for sparse signal recovery, particularly in compressed sensing. It identifies the sparsest solution to an underdetermined system by minimizing the $\ell_{1}$-norm of the signal. This approach is crucial in scenarios where data is acquired under constraints, and the signal is assumed to be sparse in certain domains. In [66], the authors observed that existing sparse channel estimation schemes for downlink CSI often deliver suboptimal reconstruction due to the use of random measurement matrices. To address this, they proposed deep unfolding-based autoencoders that optimize the measurement matrix by unfolding the basis pursuit algorithm. This customized design enhances reconstruction performance in sparse channel estimation tasks.
Sparse Bayesian learning is another effective approach for channel recovery [170]. It estimates the non-zero elements of the channel matrix while utilizing prior knowledge of sparsity. Hu et al. [67] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for multi-user MIMO systems. DDPG, a variant of deep Q-networks, is designed to handle problems in continuous action spaces [171]. Unlike traditional deep unfolding neural networks with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem. This adaptive design enhances the efficiency and performance of sparse channel recovery in varying scenarios. The channel estimation problem is formulated using an off-grid sparse Bayesian learning model. The sparse Bayesian learning algorithm is unfolded layer by layer, introducing trainable parameters such as grid points and off-grid variables. Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized mean square error performance.
V-B mmWave Massive MIMO
mmWave massive MIMO operates in the high-frequency range of 30 GHz to 300 GHz, allowing higher dimensionality and the arrangement of many antennas in a compact space due to compressed wavelengths. The BS, equipped with numerous antennas, must acquire downlink CSI. However, the large amount of feedback required creates significant overhead, posing a challenge in mmWave massive MIMO systems. Compressed sensing techniques are widely used to reduce feedback overhead by exploiting channel sparsity [172]. The limited scattering properties of mmWave propagation and the power-focusing capability of lens antenna arrays transform the spatial domain channel into a sparse beamspace representation. Consequently, the channel estimation problem becomes a sparse signal recovery task, often addressed through compressed sensing methods. However, performing beamspace channel estimation in massive MIMO systems with limited RF chains remains a challenging task.
Compressive sensing aided deep unfolding is explored in [173], where the authors apply Onsager correction for AMP [174] and construct an unfolded model similar to [18]. The network in [175] uses deep unfolding to learn beamspace channel estimates. The authors demonstrate that this algorithm performs excellently with large antenna arrays and limited RF chains. Yi et al. in [176] introduced an unfolded network called LampResNet (Learned AMP with deep residual learning). In this model, the AMP algorithm is unfolded to estimate the channel, and deep residual learning is employed to mitigate channel noise. Normalized MSE analysis shows that the proposed networks outperform LAMP [173] and achieve performance similar to LDAMP [175]. However, exploiting channel sparsity does not always yield optimal estimation performance. To address this, the authors in [85] investigate mmWave channel sparsity in the beamspace. They formulate the estimation problem as a sparse signal recovery task and improve convergence speed by unfolding the iterative PGD algorithm. In this approach, the gradient descent step size is treated as a trainable parameter.
V-C Modern Multicarrier Modulation
For frequency-selective channels, OFDM is a preferred multicarrier modulation technique. It spreads the data across multiple orthogonal subcarriers in the frequency domain, helping to mitigate the effects of frequency-selective fading. The channel is typically modeled on a per-subcarrier basis.
To enable accurate CSI feedback by transforming the channel matrix into a sparse representation, Yiyun et al. [93] proposed a neural network consisting of compression and reconstruction modules for CSI compression and feedback. They designed the reconstruction module as an unfolded ISTA network and found that higher SNR improves reconstruction performance compared to the traditional BP method. Hu et al. [94] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance CSI feedback speed and accuracy. In this approach, the conventional ISTA algorithm is unfolded, and the regularization term is learned through end-to-end training. They modeled the regularization term as a learnable, differentiable transform, further refined in [94]. By incorporating this learnable regularization module, the model better captures CSI characteristics and adapts to different levels of channel estimation error. To improve CSI feedback accuracy and efficiency in OFDM FDD-based massive MIMO systems, while reducing complexity at the user equipment, Yangyang et al. [95] proposed an unfolded ISTA network. Their study showed that the real and imaginary components of the channel differential term share similar correlations, simplifying network design. Unfolding ISTA significantly enhances the CSI encoder’s reconstruction performance.
In [45], Guo et al. proposed a two-stage low-rank CSI feedback scheme for a low-rank OFDM mmWave channel. In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA algorithm. Fast ISTA is an improved version of the conventional ISTA algorithm. A set of thresholds is derived from the input features of the network, and to reduce the uncertainty in CSI reconstruction, the iterative steps are mapped to layers of a neural network, with $L$ stages corresponding to $L$ iterations of fast ISTA. In [96], the authors explore quantized compressed CSI at the user equipment before feedback. They propose a deep unfolding-based bit-level CSI feedback method to mitigate the effects of accumulated quantization errors. Building upon ISTA, the authors use tied learned ISTA to reduce the number of learnable parameters, offering a faster and more accurate solution to the signal reduction problem. In tied learned ISTA, the trainable parameters are shared across all layers [177].
Jianqiao et al. [97] developed an unfolded inverse-free variational Bayesian learning method to reduce the high computational complexity of matrix inversion. The sparse channel recovery performance is enhanced by learning the a-priori parameters of the sparsely represented dictionary for channel estimation. In [98], based on approximate Bayesian inference and an extension of expectation consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation. In a lens-based beamspace massive MIMO system, this approach compensates for large channel attenuation in the mmWave band. The authors introduced a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network. This network unfolds the generalized expectation consistent signal recovery algorithm and formulates the wideband beamspace channel estimation problem as a compressed image recovery task. Stein’s unbiased risk estimator is used as the loss function. In [99], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation. They designed an unfolded multiple-measurement-vectors learned AMP network by jointly training the phase shift network and channel estimator, integrating the channel sparsity for improved performance.
Underwater acoustic communication is widely used for long-range transmissions, reaching up to tens of kilometers [178, 179, 180]. The main challenges in underwater communication include frequency attenuation, time-varying multipath, and large Doppler effects. Zhao et al. [47] explored integrating the MMSE structure within the layers of a DNN to create an interpretable model. This approach aims to build an equalizer framework for improving underwater communication performance.
Figure 6: Analog, digital and hybrid beamformer [181].
Cognitive radio is a wireless technology that continuously monitors the spectrum to identify unused channels, or spectrum holes, allowing for more efficient spectrum usage. In [100], a wideband signal compressed spectrum reconstruction algorithm is proposed to detect spectrum holes across a large band using a sub-Nyquist sampling rate. The authors integrate ISTA into layers of a CNN, improving the performance in both the probability of detection and the probability of false alarm for spectrum hole detection.

### v3

@@ V-A Massive MIMO
Beamforming in MIMO systems enhances signal transmission by steering signals toward the intended receivers using multiple antennas, thereby improving received SNR [140, 141]. Beamforming techniques can be broadly divided into three types: analog, digital, and hybrid [140]. Digital beamforming uses a digital precoding matrix with one RF chain per antenna, leading to high power consumption and complexity in large arrays. Analog beamforming employs phase shifters but struggles with inter-user interference. HBF [142, 143] addresses these issues by combining digital and analog methods, reducing RF chain count while maintaining performance and efficiency which is ideal for mmWave and massive MIMO in 5G and beyond.
Shlezinger et al. [28] discuss HBF optimization using classical, data-driven, and deep unfolding methods, highlighting deep unfolding’s efficiency with limited RF chains. For precoder design, the weighted MMSE (WMMSE) algorithm [144] maximizes the weighted sum rate via iterative updates involving matrix inversions and eigendecompositions. Given the nonconvexity of the power-constrained problem [145], a distributed inexact cyclic coordinate descent (CD) method is proposed for efficient locally optimal solutions.
A deep unfolded architecture, iterative algorithm induced deep unfolding network (IAIDNN) is proposed for precoder design in MU-MIMO systems in [146]. The iterative WMMSE algorithm is unfolded. To improve efficiency, WMMSE is reformulated to approximate matrix inversions. A generalized chain rule in matrix form helps visualize the gradient recurrence between adjacent layers during backpropagation. Using this rule, the gradients of trainable parameters are computed effectively. To improve convergence while avoiding matrix inversion, a downlink beamforming method using an unfolded PGD-WMMSE algorithm for multi-user multiple-input single-output (MISO) system is proposed in [147]. The authors replace matrix inversions with PGD and compute gradients in parallel. Unlike IAIDNN [146], this approach eliminates matrix inversion, making hardware implementation more practical. Additionally, it achieves a high weighted sum rate in fully loaded scenarios, where the number of users matches the number of transmit antennas at the BS.
In [146], learnable parameters are used to approximate matrix multiplication and inversion in WMMSE update rules. However, the network fails to exploit the inherent graph structure of wireless networks. This limitation is addressed in [148]. An unfolded WMMSE algorithm for transceiver design in multicell multi-user MIMO (MU-MIMO) interference channels with local CSI is proposed in [148]. The authors use graph convolutional networks (a variant of GNN) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems. Graph filters and convolutional networks approximate WMMSE-related variables, such as weighted beamformer matrices, by leveraging graph learning techniques. However, the approach in [148] lacks flexibility in learning the exact transformation of WMMSE variables. To overcome this, [149] introduces a hybrid unfolded WMMSE algorithm for transmit beamforming in MU-MIMO ad-hoc networks. This method leverages GNNs to learn functional transformations of WMMSE parameters. Compared to the classical WMMSE algorithm, the proposed network reduces the number of layers, leading to faster convergence of the optimization problem. In [150], the joint optimization of precoders for multicell MU-MIMO systems is studied. The WMMSE algorithm is unfolded to enable distributed and synchronized precoding across cooperative BSs. The methods in [146, 147] rely on centralized optimization, leading to high complexity and backhaul overhead. In contrast, due to the distributed precoding, the proposed network in [150] reduces computational cost while maintaining efficiency. An unfolded WMMSE-based NN for massive MU-MIMO under imperfect CSI is studied in [151]. The proposed network enhances performance as compared to [144], with low complexity by leveraging CSI statistics for near-optimal precoder solutions, thus accelerating convergence.
Shi et al. [152] designed a hybrid analog-digital transceiver using iterative gradient descent to obtain minimum SER (MSER) in massive MIMO systems. They unfolded the iterative gradient descent algorithm into a DNN, where step-sizes for the gradients are the learnable parameters. The authors claim that their proposed network matches the performance of the MSER-based iterative gradient descent algorithm while significantly reducing complexity.
Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity. However, these approaches can be further improved using advanced optimization frameworks like PGD or projected gradient ascent (PGA). A robust frequency-selective precoding algorithm that maximizes the expected weighted sum rate while considering per-antenna power constraints is developed for wideband MIMO-OFDM system in [153]. This frequency-selective precoding algorithm is then unfolded into a deep network, with the aid of PGD. The trainable parameters include step sizes and projections in each PGD layer. The proposed network achieves near-optimal sum rate with lower computational complexity. Zhu et al. [154] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm for MISO heterogeneous systems. The trainable set includes the step-size for the gradient ascent update and weighted sum rate coefficients used to construct the beamforming gradient direction. The proposed network generalizes well across different numbers of antennas, different numbers of BSs, and different network sizes especially without retraining. PGA is a variant of the gradient ascent algorithm. It updates the solution in the gradient direction to maximize the objective function and then projects it back onto the feasible set. Lavi et al. [155] unfolded the PGA algorithm to optimize hybrid precoding for multiuser downlink MIMO, assuming perfect CSI availability. To handle noisy CSI, they implemented the projected conceptual mirror prox algorithm [156, 157]. The step-sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach. The solution is computed within a fixed number of iterations, defined as network layers, ensuring a balance between interpretability and efficiency. In [158], Bilbao et al. proposed an efficient analog beamforming design to maximize the achievable rate of massive MIMO in-band full-duplex systems. The underlying beamforming problem is non-convex, primarily due to constant modulus constraints imposed by the use of analog phase shifters. To overcome this challenge, they developed an unfolded algorithm based on PGA combined with ZF projection onto convex sets. The convex feasible set arises due to the requirements of the ZF constraint and unit norm beamforming vectors. This approach leverages the structure of the zero-forcing conditions and the constant modulus constraints, achieving significantly faster convergence and improved performance compared to classical iterative methods. Besides PGA and PGD, CGD is also an iterative algorithm that falls under the gradient-based optimization approach. In [159], the CGD algorithm is unfolded for constant envelope precoding to reduce computational overhead. The Riemannian manifold optimization (MO)-based CGD is implemented across network layers. Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
A fast and robust precoding network for MU-MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [160]. The objective function is a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user. The iterative optimization framework is derived from the MSE uplink-downlink duality-based algorithm, which is unfolded to design a robust MSE duality network (RMSED-Net). RMSED-Net is robust to channel estimation errors and achieves sum rate close to the upper bound of its benchmark.
In [161], the interior point method is unfolded for precoding design. It leverages interference in multi-user MISO systems to minimize power while meeting signal-to-interference and noise ratio (SINR) constraints by training the proximal operators and Lagrangian multipliers. The authors show that the execution time is reduced by ${}50-70\%$ compared to the classical method. However, the network is limited to phase shift keying (PSK) modulation, requiring reformulation for higher order and QAM modulations.
A constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [162]. The proximal Jacobian ADMM is unfolded. The unfolded network reduces average transmit power and execution time by 98% compared to proximal Jacobian ADMM. The ADMM algorithm struggles with nonlinear discrete optimization due to biased alternating updates. To address this, iterative discrete estimation (IDE) is introduced in [163], ensuring unbiased updates and better convergence. Wang et al. in [163] proposed an algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference. To reduce hardware costs and power consumption in multi-user massive MIMO, He et al. [164] developed an unfolded network using the IDE-based algorithm in [163]. The unfolded network with learnable step-size and damping factor achieves lower BER across a range of SNR.
In [165] a learning-based channel semantic acquisition and beamforming solution is proposed for cell-free massive MIMO. The beamformer is constructed by unfolding successive over-relaxation-based linear beamforming algorithm which reduces estimation errors in HBF. The proposed approach reduces computational complexity and enhances robustness to imperfect CSI.
TABLE V: Summary of typical deep unfolding models for precoder design and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
Unfolded WMMSE [166] Pioneering deep unfolding model based on the WMMSE algorithm for downlink precoding in MU-MIMO systems. Achieves $10\%$ improvement in sum rate metric compared with the conventional WMMSE at SNR $=45$ dB. Advantages: The network has a few trainable parameters, leading to fast convergence. Disadvantages: Performance loss with ill-conditioned channel matrices because the matrix inversion in WMMSE is replaced by a trainable approximation.
IAIDNN [146] Unfolded WMMSE model for MU-MIMO precoding. Converges after $2.94$ seconds compared to $291.01$ seconds required in classical WMMSE and reduces WMMSE computational cost by up to $30\%$. Advantages: IAIDNN generalizes to varying system sizes (numbers of users and antennas) with minor adjustments. Disadvantages: High training computational complexity due to the increased number of trainable parameters.
PGA-Net [155] Unfolds PGA with trainable step sizes for hybrid precoding. Proposed network achieves better sum-rate than classical PGA with fewer iterations. Advantages: Fast convergence and generalize well with varying channel size, effective for wideband channels. Disadvantages: Lacks inter-user interference cancellation and involves high gradient computation complexity.
ManNet [181] Deep unfolding models for HBF (with both fully- and sub-connected architectures) based on Riemannian MO method. Over $600$ times faster and $6$ times complexity reduction compared to MO-AltMin algorithm [182] ,has lower complexity than PGA-Net [155]. Advantages: Sparsely connected network architecture, and thus, has small number of training parameters. Disadvantages: The dynamic RF chain-antenna connections has not been incorporated into the unfolding models.
V-B mmWave MIMO
mmWave MIMO systems rely on beamforming to mitigate multi-user interference while enhancing diversity and spectrum efficiency. Typically, the beam selection and digital precoding matrices are designed separately, leading to suboptimal solutions. To address this, Hu et al. [166] proposed a joint design approach combining DL and model-driven techniques. DRL optimizes beam selection, while a deep unfolding NN, based on iterative WMMSE, optimizes the digital precoding matrix. The goal is to maximize the system sum rate while satisfying beam selection constraints. The DRL and deep unfolding networks are trained sequentially, with one network’s output serving as the input to the other. Backpropagation and SGD are used for training and parameter updates. The authors verified the proposed framework outperforms benchmarks in sum-rate, complexity, and robustness while maintaining interpretability and scalability.
Lin et al. [167] studied a LMMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs. To minimize the system MSE by optimizing the precoder at each user equipment, the PGD algorithm is unfolded. The authors focused on training only the step-size in each PGD iteration, significantly reducing the training overhead. In [168], the authors combined deep generative models with unfolding techniques to achieve near-optimal HBF while minimizing feedback and computational complexity. The receiver estimates the channel using GAN, enabling efficient optimization of digital and analog precoders. The HBF problem is formulated as a bilevel optimization. The algorithm unfolds an iterative bi-level optimization problem into the layers of the CNN. There are two levels are upper and lower layer. In the upper layer digital precoder is optimized given a fixed analog precoder while in the lower layer the analog precoder is optimized given a fixed digital precoder. This alternating iterative steps are then unfolded. The trainable parameter set include the weights of the CNN that maps digital precoder to the analog precoder. The unfolded bilevel optimization algorithm achieved near optimal HBF with low feedback and computational cost.
V-C Next Generation Multiple Access
Rate-splitting multiple access (RSMA) is a flexible multiple access technology that merges the strengths of space division multiple access (SDMA) and NOMA [169]. It uses linearly precoded rate-splitting at the transmitter and SIC at the receivers. RSMA balances interference management by partially decoding some interference while treating the rest as noise, bridging the gap between SDMA (which treats all interference as noise) and NOMA (which fully decodes interference). In [170], the authors applied fractional programming with hyperplane fixed-point iteration to design the RSMA beamformer. The algorithm is unfolded into an NN, replacing its inner loop with a small-scale DNN to predict Lagrangian dual variables. Simulations show that the proposed approach achieves high performance while reducing computational complexity.